# Facilitating Multi-Role and Multi-Behavior Collaboration of Large Language Models for Online Job Seeking and Recruiting

### v1

@@ 3.6. Prompt Modification with Feedback
4. Experimental Setup
4.1. Data
In this paper, we conduct experiments on the online recruitment datasets sourced from “Boss Zhipin” [^1] https://www.zhipin.com, the largest online recruiting platform in China. To protect the privacy of candidates, all textual documents are anonymized by removing all personal identifying information. The preprocessed dataset contains 1,992 resumes (candidates) and 1,968 job descriptions (interviewers) in total. More details of dataset statistics are shown in Table 1.
4.2. Evaluation Metrics
Following the previous studies (Zhao et al., 2022b), we employ four widely used top-$k$ recommendation metrics to test the effectiveness of person-job matching in each side of candidates or jobs. These metrics include Recall (R@k), Precision (P@k), Normalized Discounted Cumulative Gain (NDCG@k), and Mean Reciprocal Rank (MRR@k). To further evaluate two-sided matching performance, we use Macro Precision, Recall, F1 scores to calculate the accuracy of successful matching between both parties.
To evaluate the quality of mock interview generation, we use BLEU (B@n) for n-gram overlap between generated utterances and candidates’ resumes. In addition to these automatic metrics, we further perform a human evaluation for a more comprehensive assessment of the mock interview generation quality, incorporating the following metrics: (1) Coherence: The extent to which the generated questions/answers are coherent with the dialogue context; (2) Relevance: The degree to which the generated questions/answers are relevant to the candidate’s resume; (3) Diversity: The level of variation in the generated questions/answers compared to previous utterances in the dialogue history.
4.3. Comparisons
To evaluate the effectiveness of the proposed framework for person-job matching, we select the following methods for comparison, including (1) NCF (He et al., 2017) leverages a collaborative filtering method to model the interaction between jobs and candidates; (2) PJFNN (Zhu et al., 2018a) employs convolutional neural networks to encode resumes and job descriptions separately and computes the matching score using cosine similarity; (3) APJFNN (Qin et al., 2018) leverages semantic representations for both resumes and job descriptions based on sequential networks along with attention hierarchically; (4) IPJF (Le et al., 2019) uses the multi-level interactions as supervision signals to indicate the preferences for candidates and jobs.
Moreover, to assess the quality of mock interview generation, we choose classical specialized pre-trained models, such as GPT-2 (Radford et al., 2019) and T5-base (Raffel et al., 2020), in the TecentPretrain (Zhao et al., 2022a) toolkit as the backbone, and fine-tune them in our training set. Inspired by the knowledge-grounded dialogue generation methods, we use K2R (Adolphs et al., 2021) as a representative to first generate a knowledge sentence based on the context, and then produce a final response.
Table 1. Statistics of the datasets used in the experiments.
Statistics Values
# of resumes 1,992
# of job descriptions 1,968
# of interview conversations 61,215
Avg. turns per session 3.19
Avg. tokens per dialog 560
Avg. tokens per resume 1,553
Avg. tokens per job description 405
# of context-response pairs 195,318
4.4. Implementation Details
In the proposed framework, we take advantage of the role-playing capabilities of LLMs, forming augmented mock interviews to enhance the evaluation process. We employ the open-sourced general LLM Nanbeige-16b-chat (Lab, 2023) as our backbone to play the roles of interviewers and candidates. We use the vLLM (Kwon et al., 2023) as the inference framework to make LLMs efficiently generate outputs. We set the hyper-parameters in LLMs as follows: temperature=0.9, top_p=0.6, repetition_penalty=1.1, and max_tokens=32,768. We use the training set to update and accumulate reflection memory, and then directly perform inference on the test set without any additional fine-tuning. [^2] Our code is available at https://github.com/p1nksnow/PJF_interview. For the implementation of baseline models, we uniformly use the default fine-tuning hyper-parameters for TecentPretrain (Zhao et al., 2022a) and K2R (Adolphs et al., 2021) and set the learning rate as 0.001.
5. Experimental Results
5.1. Overall Performance
We first discuss the performance comparison results of two-sided person-job matching. As presented in Table 2, we can see that content-based methods, such as APJFNN and PJFNN, perform less well than collaborative filtering-based NCF. This could be because the matching approaches in APJFNN and PJFNN are case-specific and highly dependent on the representation of textual documents. They potentially struggle to fully comprehend the organizational structure of resumes and job descriptions in our dataset. In contrast, NCF, which can leverage similar matching cases to infer the compatibility of a new case, exhibits slightly better performance. Moreover, the hybrid-based method IPJF uses both text representation and information from similar cases to yield improved matching results. Our proposed MockLLM achieves the best matching accuracy for jobs, candidates, and both sides. We incorporate mock interview conversations to help enrich the evaluation stage. Meanwhile, the handshake protocol in the proposed framework is suitable for two-sided matching, which supports being put into practical applications to allow interviewers to effectively screen suitable candidates and enable candidates to find appropriate positions.
Table 2. Performance comparison of two-sided person-job matching. Bold numbers indicate that we accept the hypothesis of model improvement relative to the best baseline at a significance test level of 0.01.
@@ 5.1. Overall Performance
TencentPretrain (GPT2) 13.74 8.76 5.45 3.40 2.53 1.23 0.58 0.27 2.26 2.05 2.08
TencentPretrain (T5-base) 17.70 13.53 10.58 8.52 2.52 1.48 0.89 0.59 3.21 3.34 3.15
TencentPretrain+K2R (GPT2) 14.24 8.50 4.74 2.62 2.72 1.33 0.60 0.27 2.34 1.99 2.16
TencentPretrain+K2R (T5-base) 18.36 13.35 9.92 7.68 2.96 1.70 0.99 0.62 3.45 3.37 3.23
@@ 5.2. Ablation Study
To verify the effectiveness and necessity of various components of the proposed framework, we conduct a comparison using the following ablation variants: (1) MockLLM w/o MI eliminates the connections between mock interview generation and evaluation, making the evaluation process rely solely on resumes and job descriptions; (2) MockLLM w/o RP removes the reflection behaviors and prompt modification, and only retains the two-stage forward process of person-job matching; (3) MockLLM w/o SU omits the strategy updates and consistently uses the initial strategies $m_{0}^{i}$ and $m_{0}^{c}$ for all mock interviews.
The ablation results are also shown in Table 2. The performance of MockLLM w/o MI degrades most significantly, indicating the necessity of the proposed framework that integrates the mock interview dialogue to support talent and job evaluation and the final matching decision-making process. The performance gap between MockLLM w/o RP and the full model demonstrates the effectiveness of the reflection memory generation module in facilitating the continuous refinement of interviewer and candidate behavior. MockLLM w/o SU is unable to generate targeted and valuable questions or answers, which confirms the importance of strategy updates to the framework.
@@ 5.3. Impact of Mock Interview Generation
As an innovative module of our framework, the quality of mock interview generation directly influences the matching evaluation criteria, so it also needs to be evaluated.
Automatic evaluation. As shown in the left part of Table  3, we can observe that the specially fine-tuned dialogue model is still not as effective as MockLLM in mock interview generation. After K2R uses generated outcomes as an intermediate reference to generate final utterances, the performance of the same backbone can be slightly improved in relevance to resumes and job descriptions. Our MockLLM can leverage the LLM role-playing capability to cooperate with interviewers and candidates. The multi-role and multi-behavior collaboration framework can indeed improve the quality of mock interview generation, making interviews more centered around resume content and job requirements.
Human evaluation. We employ three well-educated volunteer annotators to further evaluate the quality of questions and responses generated by the mock interviews. The evaluation is performed in a double-blind manner ensuring that the annotators do not know which output corresponds to which method. We randomly sample 100 cases from the test set and ask each annotator to rate the questions and responses generated by each method in the aforementioned metrics: Coherence, Relevance, and Diversity.
@@
The comparison results of human evaluation are listed in the right side of Table 3. Compared to baseline methods, our proposed MockLLM can generate more coherent questions and responses to align with the interview context. The improvement on the Relevance score indicates that the questions raised by MockLLM are more relevant to the candidate’s resume, making the topics discussed in the interview conversation accurate and valuable. In addition, the highest Diversity means that MockLLM can raise more questions from new directions based on existing questions, reducing the repetition of questions to ensure the quality of mock interviews.
@@
5.4. Impact of Number of Retrieved Cases
During the prompt modification process, we need to retrieve relevant cases from reflection memory and equip them as references into questioning/answering prompts. Therefore, we conduct this analysis of the suitable number of retrieved cases in our framework. We prompt LLMs to retrieve the top-$k$ relevant cases ($k=0,1,2,3$) from reflection memory. The results in Figure 5 show that incorporating the historical experience summarized in the reflection memory is truly helpful to continuously enhance the abilities of interviewers and candidates. However, retrieving more cases does not always lead to better results. In fact, including too many cases may introduce noise that affects the interview quality.
Figure 4. Comparison results before and after questioning prompt modification.
@@ 5.5. Qualitative Analysis
In addition to the above quantitative analysis, we will provide sufficient cases to illustrate the specific applications of the proposed framework in the qualitative analysis in this subsection.
@@
Analysis of prompt modification. Regarding our proposed prompt modification technique, we provide specific results before and after questioning prompt modification and give more elaborations for qualitative comparative analysis. Figure 4 shows that the initial questioning prompt without modification is a general instruction for all candidates. Compared to the initial strategy, the modified prompt expands more questioning requirements and insights in detail and incorporates relevant cases from reflection memory to draw on prior experience to help raise new valuable questions.
Figure 5. The impact of the number of retrieved cases.
@@ 6. Conclusion
In this paper, we propose MockLLM, a novel industrial framework that leverages LLM role-play capabilities to form mock interview conversations between interviewers and candidates, which serves as augmented evidence for better evaluation of person-job matching. We extend the role-playing scenario to a multi-role and multi-behavior collaboration paradigm, enabling a single LLM role to master multiple behaviors of interviewers and candidates in mock interview generation and two-sided evaluation. We utilize reflection memory generation and dynamic prompt optimization techniques to refine the behaviors of interviewers and candidates. This contributes to continuously improving the interviewer’s professionalism and the candidate’s interview experience to facilitate better person-job matching.

### v2

@@ 3.6. Strategy Modification with Feedback
Table 1. Performance comparison of two-sided person-job matching on the PJF-Base dataset. Bold numbers indicate that we accept the hypothesis of model improvement relative to the best baseline at a significance test level of 0.01.
@@
4. Experiments
4.1. Setup
Datasets. In this paper, we experiment with the online recruitment datasets sourced from “Boss Zhipin” [^1] https://www.zhipin.com, China’s largest online recruiting platform. We collect two datasets from real online platform logs for evaluation, PJF-Base and PJF-Large, with no period overlap between them. PJF-Base is carefully curated, retaining complete and high-quality resumes and job descriptions from 106 days of online data. It contains a total of 1,992 resumes and 1,968 job descriptions. To evaluate the model performance on more diverse and larger-scale data, PJF-Large has been collected from new logs over the past six months. It covers about 10k candidates and 16k job positions in total. To protect candidate privacy, all textual documents are anonymized by removing personal identification information.
Evaluation metrics. Following the previous studies (Zhao et al., 2022b), we employ four widely used top-$k$ recommendation metrics to test the effectiveness of person-job matching from both candidate and job perspectives: Recall (R@k), Precision (P@k), Normalized Discounted Cumulative Gain (NDCG@k), and Mean Reciprocal Rank (MRR@k). To further evaluate two-sided matching performance, we use Macro Precision, Recall, and F1 scores to measure the accuracy of successful matching between both parties.
To evaluate the quality of mock interview generation, we use BLEU (B@n) for n-gram overlap between generated utterances and candidates’ resumes.
Moreover, we conduct human evaluations to comprehensively assess the mock interview quality, using the following metrics: (1) Coherence: The extent to which the generated interview questions/answers are coherent with the dialogue context; (2) Relevance: The degree to which the generated interviews are relevant to the candidate’s resume; (3) Diversity: The variation in the generated interviews compared to previous dialogue history.
Comparisons. To evaluate the effectiveness of our framework for person-job matching, we compare it with the following methods: (1) NCF (He et al., 2017) leverages collaborative filtering to model the interaction between jobs and candidates; (2) PJFNN (Zhu et al., 2018a) employs convolutional neural networks to encode resumes and job descriptions separately and computes their cosine similarity scores; (3) APJFNN (Qin et al., 2018) leverages semantic representations for both resumes and job descriptions based on sequential networks and hierarchical attention; (4) IPJF (Le et al., 2019) uses the multi-level interactions as supervision signals to indicate the preferences for candidates and jobs.
For assessing the quality of mock interview generation, we fine-tune classical specialized models: GPT-2 (Radford et al., 2019) and T5-base (Raffel et al., 2020) in the TecentPretrain (Zhao et al., 2022a) toolkit on our training data for comparison. Inspired by knowledge-grounded dialogue generation methods, we use K2R (Adolphs et al., 2021) to first generate a knowledge sentence based on the context and then produce a final response. We also include a baseline named LLM-Direct, which uses the same backbone as our MockLLM but directly simulates the interview conversations without role-playing and strategy updating processes.
Table 2. Performance comparison of two-sided person-job matching on the PJF-Large dataset. Bold numbers indicate that we accept the hypothesis of model improvement relative to the best baseline at a significance test level of 0.01.
Direction Candidates Jobs Both Sides
Method NDCG@5 R@5 P@5 MRR@5 NDCG@5 R@5 P@5 MRR@5 Precision Recall F1
NCF (He et al., 2017) 0.351 0.372 0.257 0.399 0.302 0.35 0.278 0.394 0.511 0.512 0.44
APJFNN (Qin et al., 2018) 0.359 0.356 0.264 0.378 0.285 0.361 0.291 0.37 0.496 0.502 0.423
PJFNN (Zhu et al., 2018a) 0.367 0.375 0.261 0.371 0.277 0.349 0.284 0.368 0.507 0.509 0.436
IPJF (Le et al., 2019) 0.372 0.418 0.282 0.416 0.325 0.387 0.306 0.422 0.529 0.533 0.467
MockLLM w/o MI 0.324 0.343 0.227 0.358 0.284 0.325 0.258 0.353 0.522 0.531 0.503
MockLLM w/o SU 0.522 0.498 0.301 0.476 0.546 0.526 0.355 0.497 0.612 0.624 0.574
MockLLM w/o Ref 0.534 0.506 0.306 0.480 0.552 0.539 0.363 0.501 0.627 0.631 0.588
MockLLM 0.566 0.523 0.367 0.515 0.574 0.544 0.372 0.521 0.643 0.658 0.601
@@ 4.1. Setup
GPT-2 13.74 8.76 5.45 3.40 2.53 1.23 0.58 0.27 2.26 2.05 2.08
T5 17.70 13.53 10.58 8.52 2.52 1.48 0.89 0.59 3.21 3.34 3.15
GPT-2+K2R 14.24 8.50 4.74 2.62 2.72 1.33 0.60 0.27 2.34 1.99 2.16
T5+K2R 18.36 13.35 9.92 7.68 2.96 1.70 0.99 0.62 3.45 3.37 3.23
LLM-Direct 19.22 13.21 10.01 7.92 9.77 4.32 2.29 1.93 3.74 3.65 3.43
@@
4.2. Overall Performance
We conduct comprehensive comparative experiments on two benchmark datasets, PJF-Based and PJF-Large, to evaluate the effectiveness of our proposed framework in person-job matching tasks. The results on PJF-Based and PJF-Large are shown in Tables 1 and 2. Among existing methods, content-based approaches such as APJFNN and PJFNN underperform relative to collaborative filtering-based models like NCF. This performance gap can be attributed to the limitations of content-based models, which heavily rely on textual representations of resumes and job descriptions. In real-world online recruitment data, such textual inputs often exhibit considerable variability and lack consistent structural formatting. In contrast, NCF leverages similar matching cases to infer the compatibility for new cases, resulting in slightly better performance. By learning latent factors based on similar candidate-job pairs, it achieves moderately better results, though still limited in handling complex textual semantics. The hybrid-based IPJF model can further enhance performance by combining textual document encodings with information from similar historical matches. This integration helps IPJF balance generalization and specificity, yielding improved matching accuracy across evaluation metrics.
Our proposed framework, MockLLM, consistently outperforms all baselines in terms of matching accuracy for candidates, jobs, and both sides. The advantage becomes even more evident when the experiment is scaled up to the larger PJF-Large dataset, where MockLLM maintains superior performance across all matching metrics. These results highlight the robustness and scalability of our approach. Unlike traditional models, MockLLM captures both semantic and interactive dimensions of the recruitment process. MockLLM benefits from modeling textual documents into role-playing LLMs, using reflection memory to consider successful matching experiences, and leveraging handshake evaluation to improve matching satisfaction for both candidates and interviewers.
@@ 4.3. Ablation Study
To verify the effectiveness of MockLLM’s components, we conduct an ablation study with the following variants: (1) MockLLM w/o MI eliminates mock interview generation, relying solely on resumes and job descriptions for evaluation; (2) MockLLM w/o Ref removes the reflection behaviors and strategy modification, retaining only the forward process of person-job matching; (3) MockLLM w/o SU omits strategy updates and consistently uses the initial strategies $m_{0}^{i}$ and $m_{0}^{c}$ for all mock interviews. The ablation results shown in Table 1 show that MockLLM w/o MI performs the worst, underscoring the importance of integrating mock interview dialogues. The performance gap between MockLLM w/o Ref and the full model demonstrates the effectiveness of the reflection memory generation in refining interviewer and candidate behaviors. Finally, MockLLM w/o SU fails to generate targeted and valuable questions or answers, confirming the necessity of strategy updates within the framework.
@@ 4.4. Impact of Mock Interview Generation
As a key module of our framework, the quality of mock interview generation directly influences the matching evaluation criteria and needs to be carefully evaluated.
Automatic evaluation. As shown in the left part of Table  3, our MockLLM outperforms specialized dialogue models in mock interview generation. Although K2R uses generated outcomes as intermediate references to slightly improve the document relevance, it remains inferior to MockLLM. Additionally, LLM-Direct also performs less effectively than MockLLM. These results suggest that the superiority of our framework stems not only from the use of LLMs but also from the thoughtful design of essential modules such as role-playing and strategy updating.
@@
Human evaluation. We also employ three well-educated volunteers to evaluate the quality of generated interview questions and responses. This double-blind evaluation ensures annotators are unaware of which method generated which output. We randomly sample 100 cases from the test set and ask each annotator to rate the questions and responses generated by each method based on the metrics: Coherence, Relevance, and Diversity. The human evaluation results are listed on the right side of Table 3. Compared to baseline methods, our proposed MockLLM can generate more coherent questions and responses to align with the interview context. The higher Relevance score indicates that the questions raised by MockLLM are more relevant to the candidate’s resume, making the interview topics accurate and valuable. Moreover, the highest Diversity means that MockLLM can generate varied questions from new directions based on existing questions, reducing repetition and enhancing the quality of mock interviews.
4.5. Scalability Analysis
To demonstrate our framework’s broader applicability beyond the specific candidates and jobs in PJF-Base, we conduct a scalability analysis on the PJF-Large dataset. We compare our framework with the strong baseline IPJF (Le et al., 2019) across varying data sizes to evaluate the impact on their matching performance. By gradually increasing the number of candidates to be evaluated on PJF-Large, we examine how the proportion of successful matches between candidates and suitable positions changes with data size.
The results in Figure 3 show that our framework maintains consistent matching performance across different data scales, with no significant fluctuations. Notably, the NDCG score of MockLLM always outperforms the baseline IPJF by about 0.2 points. These findings indicate that our proposed method not only scales effectively with larger datasets but also maintains high matching accuracy, making it more robust and useful for real-world applications under varying data sizes.
Figure 3. The effect of varying candidate sizes.
4.6. Domain Adaption Analysis
Person-job matching tasks require adaptability to varying recruitment requirements, interview assessment focuses, and response skills across different domains. Therefore, we review all job positions in the collected data and summarize them into four categories based on the job description requirements: Testing, Operation, Development, and Others.
The comparison results in Figure 4 highlight performance variations across these domains. While IPJF performs decently in Testing, its performance drops significantly in other domains, indicating a strong dependency on the domain. In contrast, MockLLM shows consistent superiority over IPJF across all job categories, underscoring the considerable adaptability of our framework to diverse domains. MockLLM can leverage interview questions from previously matched candidates in the interviewer’s reflection memory, allowing for continuous refinement of questioning strategies. Thus, as interviewers accumulate more experience in their specific domains, their questioning strategies become increasingly specialized, enhancing the framework’s ability to adapt to specific job domains.
4.7. Fairness Analysis
Ensuring fairness in person-job matching systems is critical, particularly in sensitive domains such as recruitment, where algorithmic decisions can significantly impact individuals’ career opportunities. To protect user privacy, we anonymize all resumes prior to their integration into our framework, removing any personally identifiable information that could influence the model’s behavior in unintended ways.
To further investigate potential fairness concerns, we conduct an analysis based on the sensitive attribute gender. Specifically, we divide the test set into Male and Female subsets to compare the performance of our MockLLM with the strong baseline IPJF (Le et al., 2019). The results indicate disparities in person-job matching across genders. However, our MockLLM demonstrates reduced bias compared to the baseline, suggesting that our framework mitigates fairness issues to some extent.
Figure 4. The effect of different job domains.
Table 4. Comparison of IPJF and MockLLM by gender.
Gender Method NDCG Recall Precision MRR
Male IPJF 0.391 0.444 0.212 0.454
MockLLM 0.447 0.512 0.292 0.529
Female IPJF 0.217 0.221 0.192 0.315
MockLLM 0.385 0.468 0.273 0.440
@@ 4.7. Fairness Analysis
Figure 6. Comparison results before and after questioning strategy modification.
@@ 4.8. Qualitative Analysis
Analysis of strategy modification. Regarding our proposed strategy modification technique, we provide specific results before and after questioning strategy modification and give more elaborations for qualitative comparative analysis. Figure 6 shows that the initial questioning strategy without modification is a general instruction for all candidates. Compared to the initial strategy, the modified strategy expands more questioning requirements and insights in detail and incorporates relevant cases from reflection memory to draw on prior experience to help raise new valuable questions.
4.9. Online Evaluation
To assess the assistance of the proposed framework on online recruitment, we establish a simulated online environment identical to the real-world platform and conduct internal testing. A group of real beta users is invited to submit resumes and interact with the system in realistic scenarios. We designate the original person-job matching pipeline used on the platform as Pipeline. As a comparison, we incorporate the MockLLM framework into the system as an enhanced version named Pipeline+MockLLM, which utilizes role-playing LLMs for mock interview generation and two-sided evaluation. We conduct an A/B test to compare their performance. The results of two-sided matching metrics in Table 5 indicate a significant improvement (all $p$-values¡0.01) with the adoption of our framework, revealing that MockLLM enhances the online system’s ability to find more satisfactory job matches for candidates. This demonstrates the potential gains of MockLLM for better satisfaction for job seekers and recruiters. The system has been integrated into a simulated real-world pipeline, ensuring compliance with operational constraints and its deployment readiness. The design and results of this testing phase reflect a clear pathway to broader adoption. Currently, the inference time per case after integrating our MockLLM is about 0.39s, an acceptable increase given the performance gains. This testing phase has provided valuable insights for further optimization and sets a clear path toward full deployment.
Table 5. The online evaluation results. Bold numbers indicate significant gains with all $p$-values lower than 0.01.
Metric Pipeline Pipeline+MockLLM $p$-value
Recall 0.30 0.60 3e-14
Precision 0.40 0.45 0.008
F1 Score 0.34 0.51 2e-10
Latency 0.003s 0.39s -
@@ 5. Conclusion
This paper presents MockLLM, a novel framework leveraging LLM role-playing capabilities to form mock interview conversations, enriching evidence for more effective person-job matching. By incorporating multi-agent behavior collaboration, reflection memory generation, and dynamic strategy modification, MockLLM continuously refines the behaviors of interviewers and candidates. The framework showcases strong potential for delivering scalable and adaptable solutions that advance recruitment processes beyond the limitations of traditional methods.
Acknowledgements.
This work was supported by the Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, and Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and Public Computing Cloud, Renmin University of China, the fund for building world-class universities (disciplines) of Renmin University of China.
Appendix A Implementation Details
In the proposed framework, we take advantage of the role-playing capabilities of LLMs, forming augmented mock interviews to enhance the evaluation process. We employ the open-sourced general LLM Nanbeige-16b-chat (Lab, 2023) as our backbone to play the roles of interviewers and candidates. We use the vLLM (Kwon et al., 2023) as the inference framework to make LLMs efficiently generate outputs. We set the hyper-parameters in LLMs as follows: temperature=0.9, top_p=0.6, repetition_penalty=1.1, and max_tokens=32,768. We use the training set to update and accumulate reflection memory, and then directly perform inference on the test set without any additional fine-tuning. For the implementation of baseline models, we uniformly use the default fine-tuning hyper-parameters for TecentPretrain (Zhao et al., 2022a) and K2R (Adolphs et al., 2021) and set the learning rate as 0.001.
Appendix B Prompt Templates
Interview question raising ($f_{ques}$)
Based on the interviewer you play, ask the candidate a technical question based on the Resume, Job Description, and Interview Dialogue Context (if any). The question must be closely related to the Resume/JD Content, about project experience and related skills. The questions are not allowed to be the same as before, please be as concise as possible.
Interview response generation ($g_{resp}$)
Based on the job seeker you play, answer the interviewer’s last question based on the Resume, Job Description, and Interview Dialogue Context. The answer is not allowed to be the same as before and does not conflict with the content of the applicant’s resume, please be as concise as possible.
Interview performance evaluation ($f_{eval}$)
Based on the interviewer you play, judge whether the candidate is suitable for this position based on the Resume, Job Description, and Interview Dialogue Context History. First, pay attention to whether the Job description matches the Resume and give a matching score. Then, combine the interview performance with a comprehensive matching score.
Job position evaluation ($g_{eval}$)
Based on the job seeker you play, judge whether you can get this position based on the Resume, Job Description, and Interview Dialogue Context History. First, pay attention to whether the Job description matches your Resume and give a matching score. Then, combine your interview performance with a comprehensive matching score.